A long-distance runner can run 6 mph downhill, but her speed is cut to 2 mph when she encounters a hill. If she runs the 6 miles uphill and 6 miles downhill, what is her average speed?
If possible please show the work, i want to understand how to approach the answer. Thank you!

Answer by checkley77(12569)

D=RT
6=6T
T=6/6
T=1 HOUR FOR THE DOWN HILL RUN.
6=2T
T=6/2
T=3 HOURS FOR THE UPHILL RUN.
12=R*(3+1)
12=4R
R=12/4
R=3 MPH IS THE AVERAGE SPEED.